Question 1

In recent years, Sandhill Corporation has purchased three machines. Because of frequent employee turnover in the accounting department, a different accountant was in charge of selecting the depreciation method for each machine, and various methods have been used. Information concerning the machines is summarized in the table below.
Machine Acquired Cost Residual
Value
Useful Life
(in years)
Depreciation Method
1 Jan. 1, 2018 $104,000 $12,600 5 Straight-line
2 July 1, 2019 91,250 9,150 5 Diminishing-balance
3 Nov. 1, 2019 56,434 6,570 6 Units-of-production


For the diminishing-balance method, Sandhill Corporation uses double the straight-line rate. For the units-of-production method, total machine hours are expected to be 21,680. Actual hours of use in the first 3 years were: 2019, 420; 2020, 4,110; and 2021, 5,390.

Machine 2: Diminishing-balance depreciation

Calculation End of Year
Year Carrying Amount
Beg. of Year
Depreciation
Rate
# of
months
Depreciation
Expense
Accumulated
Depreciation
Carrying
Amount
2019 $

Entry field with incorrect answer

Entry field with correct answer

%

Entry field with incorrect answer 3/126/129/1212/12

$

Entry field with incorrect answer

$

Entry field with incorrect answer

$

Entry field with incorrect answer

2020

Entry field with incorrect answer

Entry field with incorrect answer

%

Entry field with incorrect answer 6/1212/123/129/12

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with incorrect answer

2021

Entry field with incorrect answer

Entry field with incorrect answer

%

Entry field with incorrect answer 9/1212/123/126/12

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with incorrect answer

2022

Entry field with incorrect answer

Entry field with incorrect answer

%

Entry field with incorrect answer 3/129/1212/126/12

Entry field with incorrect answer

Entry field with incorrect answer

Entry field with incorrect answer

2023

Entry field with incorrect answer

Entry field with correct answer

%

Entry field with correct answer 3/126/129/1212/12

Entry field with incorrect answer

Entry field with correct answer

Entry field with correct answer
Answer #1

1)

Machine 2: Under Diminishing Balance Method:

Depreciation Rate under Diminishing Balance method = 1 / Number of Year of service * 2

= 1/ 5 Years * 2

= 40%

Depreciation Schedule under Diminishing Balance Method:

Calculation End of Year
Year Carrying Amount of Beg. Of Year (a) Depreciation Rate (b) # of months (c ) Depreciation Expense (d = a*b*c) Accumulated Depreciation (e = d + d) Carrying Amount (a - d)
2019 $91,250 40% 6/12. $18,250 $18,250 $73,000
2020 $73,000 40% 12/12. $29,200 $47,450 $43,800
2021 $43,800 40% 12/12. $17,520 $64,970 $26,280
2022 $26,280 40% 12/12. $10,512 $75,482 $15,768
2023 $15,768 40% 12/12. $6,307.20 $81,789.20 $9,461

Note: As requested, the depreciation schedule is prepared only for Machine 2 under double declining balance method. Please comment, if you need the schedules for remaining two machines.